# Branched Covers of Open Manifolds

### v1

@@ 1. Introduction
In this paper, we do not generally require that manifolds be orientable or compact. We are interested in manifolds $M^{m}$ which have the following universal property: every mm–manifold is a branched cover of $M$. We call such manifolds universal bases. It has been known since Alexander [Ale20] that spheres are universal bases in the PL category in all dimensions. A half century later it was shown that $S^{3}$ is universal even if we restrict to 3–fold covers with connected branching sets [Hil74, Hir74, Mon74]. For $S^{4}$, one needs degree 4 to achieve universality and degree 5 suffices for the branching sets to be embedded [IP02].
@@
Once the properties of $M$ are specified, the restrictions we impose on $N$ are natural: if $M$ is open, closed or orientable, then so is $N$. When $M$ is noncompact, then since $N$ surjects to $M$ we must also have $N$ noncompact. On the other hand, when $M$ is a nonorientable universal nn–base, since we do not require covers to be orientable this means that all (open or closed, as $M$) manifolds, including orientable and nonorientable ones, admit degree–$n$ branched covering maps over $M$. However, we do not produce such a nonorientable universal base and show that one does not exist in dimension two.
We classify compact universal bases in dimensions 2 and 3 and prove that $\mathbb{R}^{2}$ and $\mathbb{R}^{3}$ are universal bases using arguments ‘that do not rely on compactification. The following is an abridged summary of our results.
@@ 1.1. Definition and local description of a branched cover
Given two (possibly noncompact, nonorientable) connected mm–dimensional PL manifolds $M$ and $N$ with empty boundary, a nondegenerate map $f:N\to M$ is a branched cover if there is a nonempty codimension-2 subpolyhedron $B$ of $M$ such that the restriction of $f$ to a map $N\setminus f^{-1}(B)$ $\rightarrow M\setminus B$ is a dd–fold covering map. We refer to $d$ as the degree, $B$ as the branch locus, and to $f^{-1}(B)$ as the branch set of $f$. If $M$ is compact, we require that $N$ is compact and hence $d$ is finite. When $M$ is noncompact, we also let $d=\aleph_{0}$ for countably infinite-sheeted covers.
@@
In the statements and proofs below, we will often focus on simple branched covering maps. A dd–fold branched covering map $f$ is said to be simple if every point $b$ in the branch set has one preimage of index 2 and (d2d-2) preimages of index 1. Equivalently, the meridian of $b$ is mapped to a transposition in the symmetric group $S_{d}$ under the map $\pi_{1}(M\backslash B)\to S_{d}$ which determines the cover. A branched covering map is generically simple [BE79, Proposition 3.3], but when possible in our constructions we will explicitly produce simple covering maps.
@@ 2. Closed universal nn–bases
We establish some restrictions on the algebraic topology of closed universal nn–bases. We then consider nonorientable universal bases, contrasting even and odd dimensions. We also prove Proposition 2.5, which states that a closed 3–manifold is a universal nn–base (for some $n$) if and only if its fundamental group is finite.
@@
For any $m$ odd, a closed nonorientable mm–manifold $M^{m}$ is not a universal nn–base for any $n\in\mathbb{N}$.
@@
Let $M^{m}$ be a nonorientable mm–dimensional manifold where $m$ is odd. Since $\chi(M)=0$ while both $b_{0}(M)=1$ and $b_{m}(M)=0$, we must have $b_{j}(M)>0$ for some $0<j<m$. Then, by Lemma 2.1, $S^{m}$ is not a finite–fold branched cover of $M$, so $M$ is not a universal nn–base. ∎
@@
The condition $b_{1}(\Sigma)=0$ rules out every other surface except $\mathbb{RP}^{2}$. We will show that there is no integer $n$ such that every closed, nonorientable surface admits a degree–$n$ branched cover over $\mathbb{RP}^{2}$. Hence, the projective plane is not a universal nn–base for any $n$.
As an aside, note that every closed surface does admit a branched cover of some degree over $\mathbb{RP}^{2}$. For orientable surfaces, this holds since they admit branched covers over $S^{2}$, which itself covers $\mathbb{RP}^{2}$. So, in particular, every orientable surface admits a degree–4 branched covering map over $\mathbb{RP}^{2}$. To see that the nonorientable surface $_{h}\mathbb{RP}^{2}$ covers $\mathbb{RP}^{2}$, we can write $_{h}\mathbb{RP}^{2}$ as $S^{2}\#_{h}\mathbb{RP}^{2}$, where the connected sum disks are made equivariant with respect to the hh–fold cyclic branched cover $S^{2}\to S^{2}$. The $\mathbb{Z}/h\mathbb{Z}$ action on this surface permutes the $\mathbb{RP}^{2}$ summands and has quotient $\mathbb{RP}^{2}$. This constructs a degree–$h$ covering map $_{h}\mathbb{RP}^{2}\to\mathbb{RP}^{2}$ with two branch points. Alternatively, given the base $\mathbb{RP}^{2}$ with two branch points $p,q$, note that $\mathbb{RP}^{2}\setminus\nu(\{p,q\})$ is homeomorphic to the boundary sum of a Möbius band and an annulus. This cover is associated with the homomorphism $\pi_{1}(\mathbb{RP}^{2}\setminus\{p,q\})\to\mathbb{Z}/h\mathbb{Z}$ given by mapping a core of the Möbius band to 0 and a core of the annulus to 1. The resulting cyclic branched cover is connected, nonorientable, and has Euler characteristic h(12)+2=2hh(1-2)+2=2-h, and hence is homeomorphic to $_{h}\mathbb{RP}^{2}$.
Returning now to the task of showing that we cannot fix a single degree for all such branched covers, let $F_{h}$ denote the nonorientable genus–$h$ surface (i.e. $F\cong\#_{h}\mathbb{RP}^{2}$) and let $f:F_{h}\to\mathbb{RP}^{2}$ be a degree–$d$ branched cover. We claim that $h\equiv d\mod 2$. By possibly perturbing $f$ while preserving degree, we may assume that $f$ is a simple branched cover  [BE79, Proposition 3.3]. The advantage of passing to a simple cover is that it makes it easy to the compute the Euler characteristic of the total space $F_{h}$. Recall that a point $x$ in the branch locus of a degree–$d$ simple cover has one preimage of index 2 and (d2)(d-2) preimages of index $1$. Let $b$ denote the number of branch points of $f$. Thus, if there are $b$ points in the branch locus of $f$ in $\mathbb{RP}^{2}$, we have
@@
Proposition 2.3 shows that the conclusion of Corollary 2.2 also holds in the case $m=2$. As far as we are aware, it is not known whether an mm–dimensional nonorientable manifold can be a universal nn–base (for some $n$) for even $m>2$.
@@ 3.3. Noncompact 3–manifolds
Note that in Lemma 3.10, the simple branched covers $f_{a},f_{b}$ may restrict to unbranched covers on some components of $A$ or $B$.
@@
We can now extend $f_{a},f_{b}$ separately over each compression body $H_{a},H_{b}$ by adding one arc to the branch set within each attached 1-handle (see Lemma 6.1 in [BE79] and also Figure 7; this is where we use that $N$ is orientable, as the 1-handle in question must be orientable for the lemma to apply). In short, this extension over a 1-handle $h\cong D^{2}\times[0,1]$ is achieved by positioning the attaching region of $h$ so that
$f_{*}((x,y)\times\{0\})=f_{*}((x,-y)\times\{1\})$
for all $(x,y)\in D^{2}$, and using the involution $(x,y)\times\{t\}\mapsto(x,-y)\times\{1-t\}$ to extend $f_{a}$ or $f_{b}$ locally over $h\cong D^{2}\times[0,1]$ (see Figures 7 and 7). We shall denote the extensions of $f_{a},f_{b}$ over $H_{a},H_{b}$ also by $f_{a},f_{b}$, respectively.
@@
Now let $\Sigma=H_{a}\cap H_{b}$ be the orientable surface along which $H_{a}$ and $H_{b}$ are glued. The maps $f_{a}$ and $f_{b}$ restrict to simple dd–fold branched covering maps $\widetilde{f}_{a},\widetilde{f}_{b}:\Sigma\rightarrow S^{2}$, which are not expected to agree. However, since the covers $\widetilde{f}_{a}$ and $\widetilde{f}_{b}$ have the same domain and degree, they are equivalent by [BE79, Theorem 3.4]. That is, by composing with automorphisms of $\Sigma$ and $S^{2}$, we may assume that the two covers agree (see also [Lür71] and [Cle73]) and we will thus denote them both by $\widetilde{f}$. Then, by [BE79, Theorem 4.1] (here we use $d\geq 3$), there are homeomorphisms $\varphi:\Sigma\rightarrow\Sigma$ and $\psi:S^{2}\rightarrow S^{2}$, both of which are isotopic to the identity function on their respective domains, which satisfy $\widetilde{f}\circ\varphi=\psi\circ\widetilde{f}$.
@@
Figure 8. A simple branched cover $f$ of an open 3–manifold NN (left) over $\mathbb{R}^{3}$ (right). In this schematic, $E_{1}$ is a dark 3-ball; $E_{2}$ has one (genus–2) boundary component; $E_{3}$ has two boundary components (one genus–2 and one genus–3). The elements of the exhaustion are indicated by the different shades. In this figure, the indicated surfaces in NN on the left map to the 2–spheres on the right at the same height. The bold surfaces in NN are the boundary components of elements $E_{1}\subset E_{2}\subset\cdots$ of an exhaustion as in Proposition 3.7. On each of these surfaces, $f$ restricts to a 3–fold simple branched cover over a 2–sphere in $\mathbb{R}^{3}$ (indicated by translating horizontally in the diagram). The nonbolded 2–spheres drawn in NN map homeomorphically onto the corresponding 2–sphere in $\mathbb{R}^{3}$.
@@ 4. Further questions
Which open mm–manifolds are universal bases?
@@
We show in Corollary 2.2, nonorientable 3–manifolds are not covered by $S^{3}$ hence are not universal nn–bases, in the most general sense possible, for any $n$. However, we can still ask:
@@
Which nonorientable 3–manifolds are branch covered by all nonorientable 3–manifolds?
@@
Is $\mathbb{RP}^{4}$ a universal nn–base for some $n$?
Note that $\mathbb{RP}^{4}$ is 8–fold covered by all closed orientable 4–manifolds (since those are 4–fold branched covers of $S^{4}$ [Pie95]), but, as far as we are aware, it is not known whether all nonorientable 4–manifolds can be realized as branched covers over $\mathbb{RP}^{4}$.
Question 4.6.
@@
    Is $S^{m}$ a universal $n_{m}$–base for all $m$? If so, can we have $n_{m}=m$?
@@
The case $n=4$ is answered affirmatively in [PZ19b]. Note that they show an open 4–manifold $M$ is a $\min\{3|\operatorname{Ends}(M)|,\aleph_{0}\}$–sheeted cover of $\mathbb{R}^{4}$, rather than a $\min\{4|\operatorname{Ends}(M)|,\aleph_{0}\}$–sheeted cover, as one might guess from $S^{4}$ being a universal 4–base but not a universal 3–base. For this reason, it is hard to predict what values of $n_{m}$ should be expected in a potential answer to Question 4.6(2).
@@
Figure 11. Top: A degree–2 branched covering of a Whitehead link complement over $S^{2}\times I$, where the latter is represented as a 3-ball (gray) with the neighborhood of an arc (white) deleted. The vertical red line is the branch set. The Whitehead manifold can be built as a union of a solid torus, $S^{1}_{0}\times D^{2}_{0},$ and a countable family of Whitehead link complements $X_{1},X_{2},\ldots$. The bold (purple) circle indicates the curve in $\partial X_{1}$ that bounds a disk in $S^{1}_{0}\times D^{2}_{0}$. The other highlighted boundary curves (blue) illustrate the gluing of $X_{j},X_{j+1}$ – the “inner” curve in $X_{j}$ is glued to the “outer” curve in $X_{j+1}$.
Bottom: We successively redraw the top right diagram in order to visualize the branch locus as a tangle in $S^{2}\times I$, where the latter is this time represented in a standard way, as a product. Note that the highlighted boundary curves are horizontal translates of each other. This is essential in concatenating these braid segments, since in the cover this means that successive circles will be identified to each other, as in the construction of the Whitehead manifold.
@@
Whitehead’s construction has been generalized by McMillan (and others), who constructed uncountably many pairwise distinct contractible open 3–manifolds [McM62]. It is shown in [MJ61, Theorem 1] that every open contractible manifold can be written as an ascending union of handlebodies, so a construction analogous to ours can be applied to all contractible open manifolds.
@@
When $M$ and $N$ are compact manifolds (possibly with boundary), one may also define a branched cover as a surjective, finite-to-one, proper, open map [BE79]. When the manifolds involved are noncompact, it is reasonable to waive the assumption that the degree is finite. One may adopt the following definition.

### v2

@@ 1. Introduction
In this paper manifolds are, with a few explicit exceptions, orientable; but they may or may not be compact. We are interested in manifolds $M^{m}$ which have the following universal property: every mm–manifold is a branched cover of $M$. We call such manifolds universal bases. It has been known since Alexander [Ale20] that spheres are universal bases in the PL category in all dimensions. A half century later it was shown that $S^{3}$ is universal even if we restrict to 3–fold covers with connected branching sets [Hil74, Hir74, Mon74]; and that assuming a manifold $M^{3}$ is a homotopy 3-sphere actually suffices to conclude that $M$ is a universal 3-base [Pie92]. For $S^{4}$, one needs degree 4 to achieve universality and degree 5 suffices for the branching sets to be embedded [IP02].
@@
Once the properties of $M$ are specified, the restrictions we impose on $N$ are natural: if $M$ is open, closed or orientable, then so is $N$. When $M$ is noncompact, then since $N$ surjects to $M$ we must also have $N$ noncompact.
On the other hand, when $M$ is non-orientable it is in principle possible that all (open or closed, as $M$) manifolds, including orientable and non-orientable ones, admit degree–$n$ branched covering maps over $M$. However, in Corollary 2.2 we prove that no compact non-orientable $n$-bases exist.
We classify compact universal bases in dimensions 2 and 3 and prove that $\mathbb{R}^{2}$ and $\mathbb{R}^{3}$ are universal bases using arguments that do not rely on compactification. Our results can be summarized as follows:
@@ 1.1. Definition and local description of a branched cover
Given two (possibly noncompact, non-orientable) connected mm–dimensional PL manifolds $M$ and $N$ with empty boundary, a nondegenerate map $f:N\to M$ is a branched cover if there is a nonempty codimension-2 subpolyhedron $B$ of $M$ such that the restriction of $f$ to a map $N\setminus f^{-1}(B)$ $\rightarrow M\setminus B$ is a dd–fold covering map. We refer to $d$ as the degree, $B$ as the branch locus, and to $f^{-1}(B)$ as the branch set of $f$. If $M$ is compact, we require that $N$ is compact and hence $d$ is finite. When $M$ is noncompact, we also let $d=\aleph_{0}$ for countably infinite-sheeted covers. If $N$ is disconnected, a map $f:N\to M$ is a branched cover if its restriction to each component of $N$ meets the above criteria.
@@
In the statements and proofs below, we will often focus on simple branched covering maps. A dd–fold branched covering map $f$ is said to be simple if every point $b$ in the branch set has one preimage of index 2 and (d2d-2) preimages of index 1. Equivalently, the meridian of $b$ is mapped to a transposition in the symmetric group $S_{d}$ under the map $\pi_{1}(M\backslash B)\to S_{d}$ which determines the cover. In dimensions 2 and 3 a branched covering map is generically simple [BE79a, Proposition 3.3 and Theorem 6.5], and simple covers form an open set in the space of branched covers between compact $m$-manifolds for all $m$ [BE79a, Proposition 3.1]. When possible in our constructions we will explicitly produce simple covering maps.
@@ 2. Closed universal nn–bases
We establish some restrictions on the algebraic topology of closed universal nn–bases. We then consider non-orientable universal bases, contrasting even and odd dimensions. We also prove Proposition 2.5, which states that a closed 3–manifold is a universal nn–base (for some $n$) if and only if its fundamental group is finite.
@@
There are no closed non-orientable universal $n$-bases for any $n\in\mathbb{N}$.
@@
Let $M$ be a closed non-orientable mm–dimensional manifold. We’ll prove the corollary separately for $m$ even and $m$ odd.
The argument for $m$ even is a generalization of one found in [BPZ25]. Let $f:N\rightarrow M$ be a dd–fold branched covering. Let $w_{1}(M)\in H^{1}(M;\mathbb{Z}_{2})$ and $w_{1}(N)\in H^{1}(N;\mathbb{Z}_{2})$ be the first Stiefel-Whitney classes of $M$ and $N$ respectively, and $M]\in H_{m}(M;\mathbb{Z}_{2})$ and $N]\in H_{m}(N;\mathbb{Z}_{2})$ the respective fundamental classes.
Note that $f^{*}(w_{1}(M))=w_{1}(N)$. The most geometric way to see this is by noting that the first Stiefel-Whitney class $w_{1}(X)\in H^{1}(X;\mathbb{Z}_{2})\cong Hom_{\mathbb{Z}_{2}}(H_{1}(X;\mathbb{Z}_{2}),\mathbb{Z}_{2})$ is characterized by the fact that if $\gamma:S^{1}\to X$ is a loop in $X$, then $w_{1}(X)([\gamma])=0$ if $\gamma$ is an orientable loop, and 1 otherwise. For any loop $\gamma:S^{1}\to N$ we may isotope $\gamma$ so that its image is disjoint from $B$ and so that $f\circ\gamma$ is an embedding. Then it is clear that $\gamma$ will be orientable if and only if $f\circ\gamma$ is, since their images are identified homeomorphically by $f$.
Then
$\langle w_{1}(N)^{m},[N]\rangle$ $=\langle f^{*}(w_{1}(M))^{m},[N]\rangle$
$=\langle w_{1}(M)^{m},f_{*}([N])\rangle$
$=\langle w_{1}(M)^{m},d[M]\rangle$
$=d\langle w_{1}(M)^{m},[M]\rangle$
where all of the evaluations are considered in $\mathbb{Z}_{2}$. If $\langle w_{1}(M)^{m},[M]\rangle=0$, then any manifold $N$ with $\langle w_{1}(N)^{m},[N]\rangle=1$ therefore cannot be a dd–fold cover (for example, $N=\mathbb{RP}^{m}$). If $\langle w_{1}(M)^{m},[M]\rangle=1$, then $\langle w_{1}(N)^{m},[N]\rangle$ must have the same parity as $d$. Since $\langle w_{1}(N)^{m},[N]\rangle$ may be 0 or 1 in even dimensions this proves that there are no non-orientable universal nn–bases when $m$ is even.
The above proof can break down in some odd dimensions. For example, all 3–manifolds $N$ bound 4–manifolds, and hence we have $\langle w_{1}(N)^{m},[N]\rangle=0$ for all such $N$. Instead, we use the following argument: when $m$ is odd, we have $\chi(M)=0$, and, since both $b_{0}(M)=1$ and $b_{m}(M)=0$, we must have $b_{j}(M)>0$ for some $0<j<m$. Then, by Lemma 2.1, $S^{m}$ is not a finite–fold branched cover of $M$, so $M$ is not a universal nn–base. ∎
@@
The condition $b_{1}(\Sigma)=0$ rules out every other surface except $\mathbb{RP}^{2}$. We will show that there is no integer $n$ such that every closed, non-orientable surface admits a degree–$n$ branched cover over $\mathbb{RP}^{2}$. Hence, the projective plane is not a universal nn–base for any $n$.
As an aside, note that every closed surface does admit a branched cover of some degree over $\mathbb{RP}^{2}$. For orientable surfaces, this holds since they admit branched covers over $S^{2}$, which itself covers $\mathbb{RP}^{2}$. So, in particular, every orientable surface admits a degree–4 branched covering map over $\mathbb{RP}^{2}$. To see that the non-orientable surface $_{h}\mathbb{RP}^{2}$ covers $\mathbb{RP}^{2}$, we can write $_{h}\mathbb{RP}^{2}$ as $S^{2}\#_{h}\mathbb{RP}^{2}$, where the connected sum disks are made equivariant with respect to the hh–fold cyclic branched cover $S^{2}\to S^{2}$. The $\mathbb{Z}/h\mathbb{Z}$ action on this surface permutes the $\mathbb{RP}^{2}$ summands and has quotient $\mathbb{RP}^{2}$. This constructs a degree–$h$ covering map $_{h}\mathbb{RP}^{2}\to\mathbb{RP}^{2}$ with two branch points. Alternatively, given the base $\mathbb{RP}^{2}$ with two branch points $p,q$, note that $\mathbb{RP}^{2}\setminus\nu(\{p,q\})$ is homeomorphic to the boundary sum of a Möbius band and an annulus. This cover is associated with the homomorphism $\pi_{1}(\mathbb{RP}^{2}\setminus\{p,q\})\to\mathbb{Z}/h\mathbb{Z}$ given by mapping a core of the Möbius band to 0 and a core of the annulus to 1. The resulting cyclic branched cover is connected, non-orientable, and has Euler characteristic h(12)+2=2hh(1-2)+2=2-h, and hence is homeomorphic to $_{h}\mathbb{RP}^{2}$.
Returning now to the task of showing that we cannot fix a single degree for all such branched covers, let $F_{h}$ denote the non-orientable genus–$h$ surface (i.e. $F\cong\#_{h}\mathbb{RP}^{2}$) and let $f:F_{h}\to\mathbb{RP}^{2}$ be a degree–$d$ branched cover. We give a direct proof of the claim, also implied by Corollary 2.2, that $h\equiv d\mod 2$. By possibly perturbing $f$ while preserving degree, we may assume that $f$ is a simple branched cover  [BE79a, Proposition 3.3]. The advantage of passing to a simple cover is that it makes it easy to the compute the Euler characteristic of the total space $F_{h}$. Recall that a point $x$ in the branch locus of a degree–$d$ simple cover has one preimage of index 2 and (d2)(d-2) preimages of index $1$. Let $b$ denote the number of branch points of $f$. Thus, if there are $b$ points in the branch locus of $f$ in $\mathbb{RP}^{2}$, we have
@@
Proposition 2.3 shows that the conclusion of Corollary 2.2 also holds in the case $m=2$. As far as we are aware, it is not known whether an mm–dimensional non-orientable manifold can be a universal nn–base (for some $n$) for even $m>2$.
@@ 3.3. Noncompact 3–manifolds
Note that in Lemma 3.10, the simple branched covers $f_{a},f_{b}$ may restrict to unbranched covers on some components of $A$ or $B$. A version of Lemma 3.10 is proven for $N$ with two connected boundary components in [BE79b, Theorem 6.2]. For completeness, we include a proof that allows for arbitrarily many boundary components in the manifold $N$.
@@
We can now extend $f_{a},f_{b}$ separately over each compression body $H_{a},H_{b}$ by adding one arc to the branch set within each attached 1-handle (see Lemma 6.1 in [BE79a] and also Figure 7; this is where we use that $N$ is orientable, as the 1-handle in question must be orientable for the lemma to apply). We briefly recall why extending $f_{a}$ over a 1-handle $h\cong D^{2}\times[0,1]$ is possible. (The argument also applies to $f_{b}$.) Without loss of generality, we may assume that that the feet of the 1-handle are mapped to the same 2-disk, $D$, in $S^{2}$, which is disjoint from the branch locus; and furthermore that
$f_{a}((x,y)\times\{0\})=f((x,-y)\times\{1\}),$
where we identified the feet of $h$, namely $D^{2}\times\{0\}$ and $D^{2}\times\{1\}$, with the disks in $A\times\{0\}$ to which they are attached.
Using the involution $(x,y)\times\{t\}\mapsto(x,-y)\times\{1-t\}$, we can extend $f_{a}$ locally over $h\cong D^{2}\times[0,1]$. In order to complete the map thus obtained to a simple cover of the desired degree, we attach a 3-ball, $D^{2}\times[0,1]$, along $D^{2}\times\{0\}$ to each additional component of pre-image $f_{a}^{-1}(D)$. We then extend the map over each such ball by a homeomorphism onto $f_{a}(h)$. See Figures 7 and 7 for an illustration in degree 3. We shall denote the extensions of $f_{a},f_{b}$ over $H_{a},H_{b}$ also by $f_{a},f_{b}$, respectively.
@@
Now let $\Sigma=H_{a}\cap H_{b}$ be the orientable surface along which $H_{a}$ and $H_{b}$ are glued. The maps $f_{a}$ and $f_{b}$ restrict to simple dd–fold branched covering maps $\widetilde{f}_{a},\widetilde{f}_{b}:\Sigma\rightarrow S^{2}$, which are not expected to agree. However, since the covers $\widetilde{f}_{a}$ and $\widetilde{f}_{b}$ have the same domain and degree, they are equivalent by [BE79a, Theorem 3.4]. That is, by composing with automorphisms of $\Sigma$ and $S^{2}$, we may assume that the two covers agree (see also [Lür71] and [Cle73]) and we will thus denote them both by $\widetilde{f}$. Then, by [BE79a, Theorem 4.1] (here we use $d\geq 3$), there are homeomorphisms $\varphi:\Sigma\rightarrow\Sigma$ and $\psi:S^{2}\rightarrow S^{2}$, both of which are isotopic to the identity function on their respective domains, which satisfy $\widetilde{f}\circ\varphi=\psi\circ\widetilde{f}$.
@@
Figure 8. A simple branched cover $f$ of an open 3–manifold NN (left) over $\mathbb{R}^{3}$ (right). In this schematic, $E_{1}$ is a dark 3-ball; $E_{2}$ has one (genus–2) boundary component; $E_{3}$ has two boundary components (one genus–2 and one genus–3). The elements of the exhaustion are indicated by the different shades. In this figure, the indicated surfaces in NN on the left map to the 2–spheres on the right at the same height. The bold surfaces in $N$ are the boundary components of elements $E_{1}\subset E_{2}\subset\cdots$ of an exhaustion as in Proposition 3.7. On each of these surfaces, $f$ restricts to a 3–fold simple branched cover over a 2–sphere in $\mathbb{R}^{3}$ (indicated by translating horizontally in the diagram). The nonbolded 2–spheres drawn in NN map homeomorphically onto the corresponding 2–sphere in $\mathbb{R}^{3}$.
@@ 4. Further questions
Which open mm–manifolds, including orientable and non-orientable ones, are universal bases?
@@
We show in Corollary 2.2 that compact non-orientable 3–manifolds are not covered by $S^{3}$ hence are not universal nn–bases, in the most general sense possible, for any $n$. However, we can still ask:
@@
Which compact non-orientable 3–manifolds are branch covered by all non-orientable 3–manifolds? Similarly, which compact non-orientable 4–manifolds are covered by all non-orientable 4–manifolds?
It was recently shown [BPZ25] that $\mathbb{RP}^{4}$ is covered by all non-orientable compact manifolds, even though it is not a universal $n$-base for any fixed $n$; and that $S^{1}\widetilde{\times}S^{3}$, the twisted $S^{3}$ bundle over $S^{1}$, is not covered by all non-orientable compact 4–manifolds.
@@
    Is $S^{m}$ a universal $n_{m}$–base for all $m$? If so, can we take $n_{m}=m$?
@@
The case $n=4$ is answered affirmatively in [PZ19b]. Note that they show an open 4–manifold $M$ is a $\min\{3|\operatorname{Ends}(M)|,\aleph_{0}\}$–sheeted cover of $\mathbb{R}^{4}$, rather than a $\min\{4|\operatorname{Ends}(M)|,\aleph_{0}\}$–sheeted cover, as one might guess from $S^{4}$ being a universal 4–base but not a universal 3–base. For this reason, it is hard to predict what values of $n_{m}$ should be expected in a potential answer to Question 4.5(2).
@@
Figure 11. Top: A degree–2 branched covering of a Whitehead link complement over $S^{2}\times I$, where the latter is represented as a 3-ball (gray) with the neighborhood of an arc (white) deleted. The vertical red line is the branch set. The Whitehead manifold can be built as a union of a solid torus, $S^{1}_{0}\times D^{2}_{0},$ and a countable family of Whitehead link complements $X_{1},X_{2},\ldots$. The bold (purple) circle indicates the curve in $\partial X_{1}$ that bounds a disk in $S^{1}_{0}\times D^{2}_{0}$. We also highlight a longitude on the “inner” boundary component. The other highlighted boundary curves (blue) illustrate the gluing of $X_{j},X_{j+1}$ – the “inner” curve in $X_{j}$ is glued to the “outer” curve in $X_{j+1}$.
Bottom: We successively redraw the top right diagram in order to visualize the branch locus as a tangle in $S^{2}\times I$, where the latter is eventually represented in a standard way, with the product structure readily apparent. Note that the highlighted (blue) curves in $S^{2}\times\partial(I)$ are horizontal translates of each other. This is essential in concatenating these braid segments, since in the cover this means that successive circles will be identified to each other, guaranteeing that the construction of the Whitehead manifold is compatible with the branched covering maps from each $X_{i}$ to its corresponding copy of $S^{2}\times I$.
@@
Remark 4.9.
We note that there is a quiet ambiguity in the construction given in our Example 4.8. We have regarded the Whitehead manifold as built by iteratively attaching copies of $X_{i}$ to $S^{1}_{0}\times D^{2}_{0}\bigcup(\cup_{j=1}^{i-1}X_{j})$, a manifold with $S^{1}\times S^{1}$ boundary. If $X_{i}$ were a solid torus, in order to determine the attachment, it would suffice to specify the image of a meridian. In our case, in order to pin down the Whitehead manifold, we technically ought to specify the images of the oriented meridian and longitude of the component of $\partial(X_{i})$ along which we perform the attachment. Even the orientation of the meridian plays a role: while $X_{i}$ admits an involution reversing the orientation of the meridian, the two choices of orientation force distinct identifications of the endpoints or the branching set, when stacking adjacent tangles in $S^{2}\times I$, in order for the gluing of each $X_{i}$ to be compatible with the covering map. Thus, if we change the orientation of the meridian on the relevant boundary component of $X_{i}$, this would result in swapping the “inside” and “outside” of the equatorial (blue) circle in $S^{2}\times I$, altering the induced identification between the tangles and thus the total branching set. Since there is not a unique choice of identifying the boundary torus so as to produce a contractible manifold, we are not too careful to remain faithful to the original construction. Specifically, if we denote $S^{3}\backslash\large(S^{1}_{0}\times D^{2}_{0}\bigcup(\cup_{j=1}^{i}X_{j})\large)=:U_{i},$ then each $U_{i}$ is an unknotted solid torus in $S^{3}$ and “the” Whitehead manifold can also be written as the union of the nested tori $S^{1}_{0}\times D^{2}_{0}\subset U_{1}\subset U_{2}\subset$.... For contractibility, it suffices that, for each $i$, the longitude of $U_{i}$ is nullhomotopic in $U_{i+1}$; this is the case for the nested union pictured. Indeed, Whitehead’s construction has been generalized far more broadly. For example, McMillan constructed uncountably many pairwise distinct contractible open 3–manifolds [McM62].
Remark 4.10.
It is shown in [MJ61, Theorem 1] that every open contractible manifold can be written as an ascending union of handlebodies. Thus, a construction analogous to ours can be applied to all contractible open manifolds.
@@
When $M$ and $N$ are compact manifolds (possibly with boundary), one may also define a branched cover as a surjective, finite-to-one, proper, open map [BE79a]. When the manifolds involved are noncompact, it is reasonable to waive the assumption that the degree is finite. One may adopt the following definition.